a) Find "\ lim_{x\to 0}\frac{\left(tanxsec^2x-x\right)}{x^3}"

 

 b) Examine whether the equation, x3- 11x +9 =0 has a real root in the interval [0,1]


 c) Check whether the following series are convergent or not (4) 

 i) "\sum _{n=1}^{\infty }\:\frac{\left(3n-1\right)}{7^n}"

(ii) "\sum _{n=1}^{\infty }\frac{\left(\:\sqrt{n^2+3}-\sqrt{n^2-3}\right)}{\sqrt{n}}\:"




1
Expert's answer
2022-05-24T15:56:33-0400

a)


"\lim\limits_{x\to 0}\dfrac{\tan x\sec^2 x-x}{x^3}=\lim\limits_{x\to 0}\dfrac{\tan x(1)^2-x}{x^3}"

"=\lim\limits_{x\to 0}\dfrac{x+\dfrac{x^3}{3}+\dfrac{2x^5}{15}+...-x}{x^3}"

"=\lim\limits_{x\to 0}\dfrac{\dfrac{x^3}{3}+\dfrac{2x^5}{15}+...}{x^3}"


"=\lim\limits_{x\to 0}(\dfrac{1}{3}+\dfrac{2x^2}{15}+...)=\dfrac{1}{3}"

b)

"f(x)=x^3-11x+9" is continuous on "[0, 1]."


"f(0)=(0)^3-11(0)+9=9>0"

"f(1)=(1)^3-11(1)+9=-1<0"

Then by the Intermediate Value Theorem there exists a number "c" in "(0, 1)" such that "f(c)=0."

Therefore the equation, "x^3- 11x +9 =0" has a real root in the interval "[0,1]" by the Intermediate Value Theorem.


c)

(i)


"\displaystyle\sum_{i=1}^{\infin}\dfrac{3n-1}{7^n}"

Use the Ratio Test


"|\dfrac{a_{n+1}}{a_n}|=|\dfrac{\dfrac{3(n+1)-1}{7^{n+1}}}{\dfrac{3n-1}{7^n}}|=\dfrac{1}{7}(\dfrac{3n+2}{3n-1})\to\dfrac{1}{7}<1"

Thus, by the Ratio Test, the given series is absolutely convergent and therefore convergent.


(ii)


"\displaystyle\sum_{i=1}^{\infin}\dfrac{\sqrt{n^2+3}-\sqrt{n^2-3}}{\sqrt{n}}"

"=\displaystyle\sum_{i=1}^{\infin}\dfrac{n^2+3-(n^2-3)}{\sqrt{n}(\sqrt{n^2+3}+\sqrt{n^2-3})}"

"=\displaystyle\sum_{i=1}^{\infin}\dfrac{6}{\sqrt{n}(\sqrt{n^2+3}+\sqrt{n^2-3})}"

Use the Limit Comparison Test

Take "a_n=\dfrac{6}{\sqrt{n}(\sqrt{n^2+3}+\sqrt{n^2-3})}, b_n=\dfrac{1}{n}"


"\lim\limits_{n\to \infin}\dfrac{a_n}{b_n}=\lim\limits_{n\to \infin}\dfrac{\dfrac{6}{\sqrt{n}(\sqrt{n^2+3}+\sqrt{n^2-3})}}{\dfrac{1}{n}}"


"=\lim\limits_{n\to \infin}\dfrac{6}{\sqrt{1+3/n^2}+\sqrt{1-3/n^2}}=3"

The harmonic series "\displaystyle\sum_{i=1}^{\infin}\dfrac{1}{n}" is divergent.

Therefore the series "\displaystyle\sum_{i=1}^{\infin}\dfrac{\sqrt{n^2+3}-\sqrt{n^2-3}}{\sqrt{n}}" is divergent by the Limit Comparison Test.